OpenStudy (anonymous):

write the standard form of the equation of the line through the given point with the given slope. through: (4,-2), slope=-1

OpenStudy (anonymous):

need help me formular of it

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Plug x1=4 y1=-2 and m=-1

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so y+2=-1(x-4)

OpenStudy (anonymous):

result is x+y=2

OpenStudy (anonymous):

ok thank you